Find a root of the following equations correct to three decimals using Newton's raphson method: x³ -5x +3=0


1
Expert's answer
2021-06-08T04:25:11-0400
"x^3-5x+3=0"

"f(x)=x^3-5x+3"

"f'(x)=3x^2-5"

"x_{n+1}=x_n-\dfrac{f(x_n)}{f'(x_n)}"

"x_0=-2"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 step & x_0 & f(x_0) & x_1 & f(x_1) \\ \hline
 1 & -2.0000 & 5 & -2.7143 & -3.4257 \\
 \hdashline
 2 & -2.7143 & -3.4257 & -2.5140 & -0.3187 \\
 \hdashline
 3 & -2.5140 & -0.3187 & -2.4912 & -0.0039 \\
 \hdashline
 4 & -2.4912 & -0.0039 & -2.4909 & 0.0000 \\
 \hdashline
5 & -2.4909& -0.0000 & -2.4909 & 0.0000 \\
 \hdashline
\end{array}"



"x_{I}=-2.491"



"x_0=1"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 step & x_0 & f(x_0) & x_1 & f(x_1) \\ \hline
 1 & 1.0000 & -1.0000 & 0.5000 & 0.6250 \\
 \hdashline
 2 & 0.5000 & 0.6250 & 0.6471 & 0.0356 \\
 \hdashline
 3 & 0.6471 & 0.0356 & 0.6566 & 0.0002\\
 \hdashline
 4 & 0.6566 & 0.0002 & 0.6566 & 0.0002 \\
 \hdashline

\end{array}"



"x_{II}=0.657"




"x_0=2"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 step & x_0 & f(x_0) & x_1 & f(x_1) \\ \hline
 1 & 2.0000 & 1.0000 & 1.8571 & 0.1195 \\
 \hdashline
 2 & 1.8571 & 0.1195 & 1.8348 & 0.0028 \\
 \hdashline
 3 & 1.8348 & 0.0028 & 1.8342 & 0.0000\\
 \hdashline
 4 & 1.8342 & 0.0000 & 1.8342 & 0.0000 \\
 \hdashline

\end{array}"



"x_{III}=1.834"